A student measures the initial temperature of an object as 22.75 oC and the final temperature as 55.87 oC. What is the change in temperature in units of Kelvin?
Answer #1

First of all the temperatures are given in degree celsius, therefore we will convert the temperature into kelvin as the answer has been asked in units of kelvin.

FORMULA FOR CONVERSION OF TEMPERATURE FROM DEGREE CELSIUS TO KELVIN

The temperature in degree celsius + 273.15

Therefore initial temperature in Kelvin

22.75 + 273.15 = 295.9

Final temperature in Kelvin

55.87 + 273.15 = 329.02

Change in temperature = Final temperature - Initial temperature

= 329.02 - 295.9

= 33.12 kelvin